twenty years ago, a mother was 4 times as old as her son 4 years hence,she will be twicw as old as her son .what are their present ages
Answer by macston(5194)

.
M=mother's age; S=son's age
.
M-20=4(S-20)
M=4(S-20)+20
M=4S-80+20
M=4S-60
.
M+4=2(S+4)
M=2(S+4)-4
M=2S+8-4
M=2S+4
.
4S-60=2S+4
2S-60=4
2S=64
S=32
.
M=4S-60=4(32)-60=128=60=68
.
ANSWER: Son is 32 and mother is 68.
.
CHECK:
M+4=2(S+4)
68+4=2(32+4)
72=2(36)
72=72
.